An object is placed 6 cm from a convex lens of focal length 18 cm. what is the lateral magnification of the object.
Answer #1

Lens formula for Convex lens is given as follows :-

(1 / v) - (1/u) = (1/f)

so ( 1/ v ) = ( 1 / u ) + ( 1 / f )

1 / v = ( 1/6 ) + ( 1/18 )

v = 4.5 cm

so Lateral magnification = v / u

LM = 4.5 / 6 = 0.75 <<<<<answer